A current of 350 mA flows through a gold wire with a diameter of 1.0 mm and a length of 25 cm. There are 5.90 x 1028 free electrons/m3in a gold. 


A) What is the current density in the wire?

B) What is the magnitude of the drift velocity of the free electrons in the wire?




1
Expert's answer
2021-03-14T19:21:46-0400

(a) By the definition of the current density, we have:


"J=\dfrac{I}{A}=\dfrac{4I}{\pi d^2}=\dfrac{4\cdot350\cdot10^{-3}\ A}{\pi\cdot(1.0\cdot10^{-3}\ m^2)^2}=4.46\cdot10^{-5}\ \dfrac{A}{m^2}."

(b) The magnitude of the drift velocity of the free electrons in the wire can be found as follows:


"v=\dfrac{J}{n|e|}=\dfrac{4.46\cdot10^{-5}\ \dfrac{A}{m^2}}{5.9\cdot10^{28}\ \dfrac{el}{m^3}\cdot|-1.6\cdot10^{-19}\ C|}=4.72\cdot10^{-15}\ \dfrac{m}{s}."